OpenStudy (anonymous):

Triangle ABC Has Vertics At A(3,0) B(9,-5) And C(7,-8). Find The Length of AC Simplest Radical Form.

OpenStudy (anonymous):

4 times the square root 5

OpenStudy (wolf1728):

I haven't made the calculations yet but here is the graphic

OpenStudy (wolf1728):

AC² = 8² + 4² AC² = 64 + 16 AC = sqrt(80) AC = 8.94427191 or as the previous answer stated AC = 4 * sqrt(5)